Based on historical data, your manager believes that 36% of the company's orders come from first-time customers. A random sample of 201 orders will be used to estimate the proportion of first-time-customers.

What is the probability that the sample proportion is between 0.32 and 0.45?
Answer #1

Solution:

Given that,

P = 36% = 0.36

1 - P = 1 - 0.36 = 0.64

n = 201

\mu\hat p= 0.36

\sigma\hat p = \sqrt (P ( 1 - P ) / n )

= \sqrt 0.36 * 0.64 / 201

= 0.0338

P ( 0.32 < \hat p < 0.45 )

( 0.32 - 0.36 / 0.0338) < ( \hat p -   \mu\hat p / \sigma \hat p ) < ( 0.45 - 0.36 / 0.0338 )

p ( - 0.04 /0.0338 < z < 0.09 / 0.0338 )

p (-1.18 < z < 2.66 )

p ( z < 2.66) - p ( z < -1.18 )

Using z table

= 0.9961 - 0.1190

= 0.8771

Probability = 0.8771